OpenStudy (anonymous):

The function S(n) is defined below. Which of the following are equal to S(10)? Check all that apply. S(n) = 5n - 3 A. S(9) + 5 B. 53 C. 47 D. S(7) + S(3)

OpenStudy (godlovesme):

s(n)=5n-3 s(10)= 5(10) -3

OpenStudy (idku):

wierd usually, S stands for a sum...

OpenStudy (anonymous):

it would be C (47)

OpenStudy (anonymous):

Hi tallan Firstly, we are told: S(n) =5n-3 if n=10, then: S(10) this means substitute 10 in the right hand side everywhere you see n i.e S(10) = 5(10) -3 S(10) = 50 -3 S(10) = 47

OpenStudy (anonymous):

Ok I understand that now.

OpenStudy (anonymous):

cool!